Mr. Abbot left Farmersville in a plane at noon to travel to Exeter. Mr. Baker left Exeter in his automobile at 2pm to travel to Farmersville. It is 400 mi from Exeter to Farmersville. If the sum of their speeds was 120mph, and if they crossed paths at 4pm, find the speed of each.
x= speed of plane
y= speed if automobile
The system is
x+y=400
4x+4y=120

Found 2 solutions by mananth, slitco:
Answer by mananth(16946)


x= speed of plane
y= speed if automobile
x+y=120
4x+2y=400 ( he leaves at 2.00 pm)
..
4x+4y-4x-2y=480-400
2y=80
y=40 mph speed of car.
x+y=120
so x 80 mph.

Answer by slitco(2)

p = speed of plane
a = speed of automobile
------------------------
p + a = 120
but were gonna make this "a = -p + 120" for the sake of the problem bc that's what works lol.
then the other equation will be 4p + 2a = 400
-------------------------------------------------------------------
4p + 2(-p + 120) = 400
4p - 2p + 240 = 400
2p + 240 = 400
2p = 160
p = 80 mph (plane)
then plug 80 into -p in the other equation...which is a = -p + 120
and that gives you...
a = 40 mph (automobile)
:) <3